# Dark matter from phase transition generated PBH evaporation with gravitational waves signatures

### v1

@@ I Introduction
Primordial black holes (PBH), originally proposed in Zeldovich:1967lct and later by Hawking Hawking:1974rv ; Hawking:1975vcx can have several interesting cosmological consequences Chapline:1975ojl ; Carr:1976zz . A recent review of PBH can be found in Carr:2020gox . Depending upon the initial mass $(m_{\rm in})$ and initial energy fraction $(\beta)$, PBH can face different astrophysical and cosmological constraints. Among them, the ultra-light PBH mass window is relatively less constrained as such PBH evaporate by emitting Hawking radiation Hawking:1974rv ; Hawking:1975vcx before the big bang nucleosynthesis (BBN) epoch. Comparing the evaporation temperature with the BBN temperature $T_{\text{ev}}>T_{\text{BBN}}\simeq 4$ MeV leads to the upper bound on this ultra-light PBH mass window. On the other hand, a lower bound can be obtained by using the upper limit on the scale of inflation from cosmic microwave background (CMB) measurements. Using these BBN and CMB bounds simultaneously leads to the allowed initial mass for ultra-light PBH that reads $0.1\,\text{g}\lesssim m_{\text{in}}\lesssim 3.4\times 10^{8}\,\text{g}$. As mentioned above, the range of PBH masses in this window is relatively unconstrained Carr:2020gox . It should be noted that the upper bound is obtained only for PBH which evaporates on a timescale smaller than the age of the universe and have a large initial fraction $\beta$ to dominate the energy density at some epoch. Such evaporation of ultra-light PBH can also lead to the production of dark matter (DM) in the universe Gondolo:2020uqv ; Bernal:2020bjf ; Green:1999yh ; Khlopov:2004tn ; Dai:2009hx ; Allahverdi:2017sks ; Lennon:2017tqq ; Hooper:2019gtx ; Sandick:2021gew ; Fujita:2014hha ; Datta:2020bht ; JyotiDas:2021shi ; Barman:2021ost ; Barman:2022gjo ; Barman:2022pdo ; Cheek:2021odj . As suggested by observations in astrophysics and cosmology based experiments, approximately $26\%$ of the present universe is made up of DM, the particle origin of which is not yet known. If DM has only gravitational interactions as observations indicate, the conventional thermal freeze-out mechanism fails and PBH evaporation provides a very good alternative.
@@
While DM with only gravitational interactions may not have much detection prospects, the ultra-light PBH and their formation mechanisms in the early universe can lead to promising detection prospects. In spite of this relatively less constrained mass window, it is still possible to probe an ultra-light PBH dominated epoch in the early universe via detection of stochastic gravitational waves (GW) background. Interestingly, detection of stochastic GW can also probe the formation mechanism of PBH in the early universe. In the early universe, PBH can be formed in a variety of ways like, from inflationary perturbations Hawking:1971ei ; Carr:1974nx ; Wang:2019kaf ; Byrnes:2021jka ; Braglia:2022phb , first-order phase transition (FOPT) Crawford:1982yz ; Hawking:1982ga ; Moss:1994iq ; Kodama:1982sf ; Baker:2021nyl ; Kawana:2021tde ; Huang:2022him ; Hashino:2021qoq ; Liu:2021svg ; Gouttenoire:2023naa ; Lewicki:2023ioy , the collapse of topological defects Hawking:1987bn ; Deng:2016vzb etc. Such mechanisms often come up with their own unique GW spectrum. On the other hand, PBH can lead to generation of GW in a variety of ways. The Hawking evaporation of PBH itself can produce gravitons which might constitute an ultra-high frequency GW spectrum Anantua:2008am . PBH can also lead to GW emission by forming successive mergers Zagorac:2019ekv . In addition, the scalar perturbations leading to the formation of PBH can induce GW at second-order Saito:2008jc , which can be further enhanced during the PBH evaporation Inomata:2020lmk . Finally, the inhomogeneity in the distribution of PBH may also induce GW at second order, as recently studied in Papanikolaou:2020qtd ; Domenech:2020ssp ; Domenech:2021wkk . In the present work, we consider the last possibility which, for ultra-light PBH mass window, can lead to stochastic GW with peak frequencies in the ballpark of ongoing and planned future experiments.
@@
Motivated by this, we study the possibility of simultaneously probing the signatures of formation mechanism and PBH via detection of stochastic GW background with non-trivial implications for dark matter produced solely from PBH evaporation. Recently, FOPT origin of heavy PBH was studied by the authors of Gehrman:2023esa ; Xie:2023cwi ; Banerjee:2023brn ; Baldes:2023rqv ; Banerjee:2023vst . Here, we consider the ultra-light PBH mass window, not considered in earlier works and study the possibility of forming them in FOPT. The combined GW spectrum arising out of the FOPT and density perturbations of ultra-light PBH consists of a unique doubly-peaked feature. For the PBH mass range of interest, the peak in GW spectrum due to FOPT occurs at a frequency higher than the one due to PBH density perturbations. Depending upon the model parameters, the peak frequencies and corresponding blue and red tilted parts of the GW spectrum remain within the sensitivities of several planned experiments like DECIGO Kawamura:2006up , BBO Yagi:2011wg , ET Punturo_2010 , CE LIGOScientific:2016wof , UDECIGO (UDECIGO-corr) Sato:2017dkf ; Ishikawa:2020hlo etc. While doubly peaked GW spectrum arises in a scenario with PBH domination, the DM parameter space is typically restricted to the superheavy ballpark. In sub-dominant PBH scenario, DM mass range gets much wider whereas the GW spectrum generated by FOPT comes within the sensitivities of several future experiments. Depending upon the specific scenario, DM masses from GeV to superheavy regime and PBH mass as low as a few kg remains within sensitivities of such GW experiments. It is worth mentioning that in most of the earlier works studying DM production from PBH evaporation, the formation mechanism of the latter is not included. The inclusion of the PBH formation mechanism not only leads to a tighter parameter space but also to complementary detection prospects.
@@
This paper is organised as follows. In section II, we outline the particle physics setup responsible for FOPT followed by discussion of PBH formation in section III. In section IV and section V, we discuss the stochastic GW generated by FOPT and PBH density fluctuations respectively. In section VI, we discuss the combined GW spectrum arising from FOPT and PBH followed by implications for DM phenomenology in section VII. We finally conclude in section VIII.
@@ II The setup
Among different scenarios relating FOPT to PBH formation Crawford:1982yz ; Hawking:1982ga ; Moss:1994iq ; Kodama:1982sf ; Baker:2021nyl ; Kawana:2021tde ; Huang:2022him ; Hashino:2021qoq ; Liu:2021svg ; Gouttenoire:2023naa ; Lewicki:2023ioy , we consider the false vacuum collapse as the origin of PBH formation. In particular, we consider the collapsing Fermi-ball scenario Kawana:2021tde ; Huang:2022him ; Marfatia:2021hcp ; Tseng:2022jta ; Lu:2022jnp ; Marfatia:2022jiz ; Xie:2023cwi . In this scenario, a dark sector fermion $\chi$ with an asymmetry between $\chi$ and $\bar{\chi}$ leads to the formation of Fermi-balls after getting trapped in the false vacuum. Due to the strong Yukawa force among the dark fermions, such Fermi-balls can collapse into PBH. As the dark sector is asymmetric [^1] Here, we remain agnostic about the origin of such an asymmetry. Example of a UV complete model can be found in Huang:2022him . The GW phenomenology discussed here is generic and independent of such UV completions. and carries a conserved charge, we consider a global dark $U(1)_{D}$ symmetry which remains unbroken as the Universe goes through a FOPT. The dark sector fermion $\chi$ has a charge $q_{\chi}$ under this $U(1)_{D}$ global symmetry while all other fields are neutral. We consider the FOPT to be driven by a singlet scalar $\Phi$. The dark sector fermion $\chi$ is coupled to $\Phi$ through Yukawa interaction. An additional singlet scalar $\Phi^{\prime}$ and doublet scalar $H_{2}$ are also considered such that we have a strong FOPT while keeping all dimensionless quartic couplings in perturbative ballpark. These scalars are also neutral under $U(1)_{D}$ global symmetry and their couplings with $\Phi$ are the only relevant ones for our analysis.
@@ III PBH formation from Fermi-ball
As the FOPT proceeds, the true vacuum expands to cover the whole Universe, while the false vacuum gradually shrinks to a smaller region. At a later stage, the false vacuum region contracted into smaller disconnected volumes, which again split to negligible size at percolation temperature Ellis:2018mja . Within these final pockets of false vacuum, some pre-existing asymmetry of $\chi$ fermions can get trapped and the symmetric part annihilates via $\chi\chi\rightarrow$S and $\chi\chi\rightarrow$SS, to final state particles S (either SM particles or $\phi$) through decay or annihilation. Now, the trapped $\chi$’s create a degeneracy pressure. If such a pressure is able to balance the vacuum pressure, these $\chi$ remnants form a bound state which is known as Fermi-ball Lee:1986tr ; Hong:2020est ; Kawana:2021tde . These Fermi-balls formed at $T_{*}$ defined by $p(T_{*})$=0.29. The energy of a Fermi-ball can be written as
@@
$E_{Y}\simeq-\frac{3g_{\chi}^{2}}{8\pi}\frac{Q_{\rm FB}^{3}}{R_{\rm FB}}\left(\frac{L_{\phi}}{R_{\rm FB}}\right)^{2}.$ (13)
@@
$\beta\gtrsim\beta_{\text{crit}}\simeq\frac{T_{\rm evap}}{T_{n}}\,,$ (16)
@@
$T_{n}=1.6\times 10^{18}~{}\beta_{\rm PBH}\left(\frac{100}{g_{*}}\right)^{1/4}\left(\frac{M_{\rm in}}{1\hskip 2.84544pt\rm g}\right)^{-1/2}\left(\frac{\eta_{\chi}}{10^{-3}}\right)^{-1}\frac{1}{\alpha_{*}^{1/4}}\hskip 2.84544pt\text{GeV}\,,$ (17)
$\beta/\textbf{H}_{*}=1.2\times 10^{-2}\alpha^{1/4}_{*}\left(\frac{\eta_{\chi}}{10^{-3}}\right)v_{w}\left(\frac{g_{*}}{100}\right)^{1/12}\beta_{\text{PBH}}^{-2/3}\,.$ (18)
@@ V GW from PBH density fluctuations
While there are several ways in which PBH can give rise to GW Anantua:2008am ; Zagorac:2019ekv ; Saito:2008jc , for our purpose of ultralight PBH, the GW spectrum generated by most of these mechanisms lies at very high frequencies. Hence, we rather focus on GW generated because of the inhomogeneity in the PBH distribution after they are formed Papanikolaou:2020qtd ; Domenech:2020ssp ; Domenech:2021wkk . Such a spectrum is independent of the formation mechanism of PBH, and is within reach of near-future GW detectors.
@@
where $k_{\rm UV}$ corresponds to the comoving wavenumber corresponding to the mean separation between PBH, whereas $k_{\rm eq}$, $k_{\rm eva}$ indicates the comoving wavenumber of the modes entering the horizon when PBH begins to dominate and when they evaporate respectively. The ratio of these wavenumbers appearing in Eq. (31) depends on the PBH mass and abundance, and for PBH formed through the collapse of Fermi-balls, we find the amplitude to be
@@
Note that $T_{n}$ in turn depends on the PBH parameters and the asymmetry $\eta_{\chi}$ (cf. Eq. (17)).
@@
$\beta_{\rm PBH}\lesssim\beta_{\rm max}\simeq 1.8\times 10^{-4}\left(\frac{M_{\rm in}}{10^{4}g}\right)^{-17/32}\left(\frac{\alpha}{0.1}\right)^{3/16}\left(\frac{\beta/H}{10}\right)^{3/4}\,.$ (33)

### v2

@@ I Introduction
Primordial black holes (PBH), originally proposed in Zeldovich:1967lct and later by Hawking Hawking:1974rv ; Hawking:1975vcx can have several interesting cosmological consequences Chapline:1975ojl ; Carr:1976zz . A recent review of PBH can be found in Carr:2020gox . Depending upon the initial mass $(m_{\rm in})$ and initial energy fraction $(\beta)$, PBH can face different astrophysical and cosmological constraints. Among them, the ultra-light PBH mass window is relatively less constrained as such PBH evaporate by emitting Hawking radiation Hawking:1974rv ; Hawking:1975vcx before the big bang nucleosynthesis (BBN) epoch. Comparing the evaporation temperature with the BBN temperature $T_{\text{ev}}>T_{\text{BBN}}\simeq 4$ MeV leads to the upper bound on this ultra-light PBH mass window. On the other hand, a lower bound can be obtained by using the upper limit on the scale of inflation from cosmic microwave background (CMB) measurements. Using these BBN and CMB bounds simultaneously leads to the allowed initial mass for ultra-light PBH that reads $0.1\,\text{g}\lesssim m_{\text{in}}\lesssim 3.4\times 10^{8}\,\text{g}$. As mentioned above, the range of PBH masses in this window is relatively unconstrained Carr:2020gox . It should be noted that the upper bound is obtained only for PBH which evaporates on a timescale smaller than the age of the universe and have a large initial fraction $\beta$ to dominate the energy density at some epoch. Such evaporation of ultra-light PBH can also lead to the production of dark matter (DM) in the universe Gondolo:2020uqv ; Bernal:2020bjf ; Green:1999yh ; Khlopov:2004tn ; Dai:2009hx ; Allahverdi:2017sks ; Lennon:2017tqq ; Hooper:2019gtx ; Sandick:2021gew ; Fujita:2014hha ; Datta:2020bht ; JyotiDas:2021shi ; Barman:2021ost ; Barman:2022gjo ; Barman:2022pdo ; Cheek:2021odj ; Chaudhuri:2023aiv . As suggested by observations in astrophysics and cosmology based experiments, approximately $26\%$ of the present universe is made up of DM, the particle origin of which is not yet known. If DM has only gravitational interactions as observations indicate, the conventional thermal freeze-out mechanism fails and PBH evaporation provides a very good alternative.
@@
While DM with only gravitational interactions may not have much detection prospects, the ultra-light PBH and their formation mechanisms in the early universe can lead to promising detection prospects. In spite of this relatively less constrained mass window, it is still possible to probe an ultra-light PBH dominated epoch in the early universe via detection of stochastic gravitational waves (GW) background  [^1] See Ref. Roshan:2024qnv for a recent review on stochastic gravitational wave background.. Interestingly, detection of stochastic GW can also probe the formation mechanism of PBH in the early universe. In the early universe, PBH can be formed in a variety of ways like, from inflationary perturbations Hawking:1971ei ; Carr:1974nx ; Wang:2019kaf ; Byrnes:2021jka ; Braglia:2022phb , first-order phase transition (FOPT) Crawford:1982yz ; Hawking:1982ga ; Moss:1994iq ; Kodama:1982sf ; Baker:2021nyl ; Kawana:2021tde ; Huang:2022him ; Hashino:2021qoq ; Liu:2021svg ; Gouttenoire:2023naa ; Lewicki:2023ioy ; Gehrman:2023qjn ; Kim:2023ixo , the collapse of topological defects Hawking:1987bn ; Deng:2016vzb etc. Such mechanisms often come up with their own unique GW spectrum. On the other hand, PBH can lead to generation of GW in a variety of ways. The Hawking evaporation of PBH itself can produce gravitons which might constitute an ultra-high frequency GW spectrum Anantua:2008am . PBH can also lead to GW emission by forming successive mergers Zagorac:2019ekv . In addition, the scalar perturbations leading to the formation of PBH can induce GW at second-order Saito:2008jc , which can be further enhanced during the PBH evaporation Inomata:2020lmk . Finally, the inhomogeneity in the distribution of PBH may also induce GW at second order, as recently studied in Papanikolaou:2020qtd ; Domenech:2020ssp ; Domenech:2021wkk . In the present work, we consider the last possibility which, for ultra-light PBH mass window, can lead to stochastic GW with peak frequencies in the ballpark of ongoing and planned future experiments.
@@
Motivated by this, we study the possibility of simultaneously probing the signatures of formation mechanism and PBH via detection of stochastic GW background with non-trivial implications for dark matter produced solely from PBH evaporation. Recently, FOPT origin of heavy PBH was studied by the authors of Gehrman:2023esa ; Xie:2023cwi ; Banerjee:2023brn ; Baldes:2023rqv ; Banerjee:2023vst ; Gouttenoire:2023pxh . Here, we consider the ultra-light PBH mass window, not considered in earlier works and study the possibility of forming them in FOPT. The combined GW spectrum arising out of the FOPT and density perturbations of ultra-light PBH consists of a unique doubly-peaked feature. For the PBH mass range of interest, the peak in GW spectrum due to FOPT occurs at a frequency higher than the one due to PBH density perturbations. Depending upon the model parameters, the peak frequencies and corresponding blue and red tilted parts of the GW spectrum remain within the sensitivities of several planned experiments like DECIGO Kawamura:2006up , BBO Yagi:2011wg , ET Punturo_2010 , CE LIGOScientific:2016wof , UDECIGO (UDECIGO-corr) Sato:2017dkf ; Ishikawa:2020hlo etc. While doubly peaked GW spectrum arises in a scenario with PBH domination, the DM parameter space is typically restricted to the superheavy ballpark. In sub-dominant PBH scenario, DM mass range gets much wider whereas the GW spectrum generated by FOPT comes within the sensitivities of several future experiments. Depending upon the specific scenario, DM masses from GeV to superheavy regime and PBH mass as low as a few kg remains within sensitivities of such GW experiments. It is worth mentioning that in most of the earlier works studying DM production from PBH evaporation, the formation mechanism of the latter is not included. The inclusion of the PBH formation mechanism not only leads to a tighter parameter space but also to complementary detection prospects.
@@
This paper is organised as follows. In section II, we outline the particle physics setup responsible for FOPT followed by discussion of PBH formation in section III. In section IV and section V, we discuss the stochastic GW generated by FOPT and PBH density fluctuations respectively. In section VI, we discuss the combined GW spectrum arising from FOPT and PBH followed by implications for DM phenomenology in section VII. We briefly comment on other FOPT origin of PBH not discussed in this work and the corresponding GW signatures in section VIII. We finally conclude in section IX.
@@ II The setup
Among different scenarios relating FOPT to PBH formation Crawford:1982yz ; Hawking:1982ga ; Moss:1994iq ; Kodama:1982sf ; Baker:2021nyl ; Kawana:2021tde ; Huang:2022him ; Hashino:2021qoq ; Liu:2021svg ; Gouttenoire:2023naa ; Lewicki:2023ioy ; Gehrman:2023qjn ; Kim:2023ixo , we consider the false vacuum collapse as the origin of PBH formation. In particular, we consider the collapsing Fermi-ball scenario Kawana:2021tde ; Huang:2022him ; Marfatia:2021hcp ; Tseng:2022jta ; Lu:2022jnp ; Marfatia:2022jiz ; Xie:2023cwi ; Gehrman:2023qjn ; Kim:2023ixo . In this scenario, a dark sector fermion $\chi$ with an asymmetry between $\chi$ and $\bar{\chi}$ leads to the formation of Fermi-balls after getting trapped in the false vacuum. Due to the strong Yukawa force among the dark fermions, such Fermi-balls can collapse into PBH. As the dark sector is asymmetric [^2] Here, we remain agnostic about the origin of such an asymmetry. Example of a UV complete model can be found in Huang:2022him . The GW phenomenology discussed here is generic and independent of such UV completions. and carries a conserved charge, we consider a global dark $U(1)_{D}$ symmetry which remains unbroken as the Universe goes through a FOPT. The dark sector fermion $\chi$ has a charge $q_{\chi}$ under this $U(1)_{D}$ global symmetry while all other fields are neutral. We consider the FOPT to be driven by a singlet scalar $\Phi$. The dark sector fermion $\chi$ is coupled to $\Phi$ through Yukawa interaction. Since fermion and bosons contribute with opposite signs to the effective potential, and fermion $\chi$ has large Yukawa coupling with $\Phi$, additional bosonic degrees of freedom are required to have an effective potential consistent with a strong FOPT while keeping all dimensionless couplings within perturbative limits. We consider an additional singlet scalar $\Phi^{\prime}$ and doublet scalar $H_{2}$ for this purpose. However, this choice is arbitrary. For example, if we embed this setup within a new gauge symmetry under which $\Phi$ transforms non-trivially, such extra scalar fields are no longer required. These additional scalars are also neutral under $U(1)_{D}$ global symmetry and their couplings with $\Phi$ are the only relevant ones for our analysis.
@@ III PBH formation from Fermi-ball
As the FOPT proceeds, the true vacuum expands to cover the whole Universe, while the false vacuum gradually shrinks to a smaller region. At a later stage, the false vacuum region contracted into smaller disconnected volumes, which again split to negligible size at percolation temperature Ellis:2018mja . Within these final pockets of false vacuum, some pre-existing asymmetry of $\chi$ fermions can get trapped and the symmetric part annihilates via $\chi\chi\rightarrow$S and $\chi\chi\rightarrow$SS, to final state particles S (either SM particles or $\phi$) through decay or annihilation. Now, the trapped $\chi$’s create a degeneracy pressure. If such a pressure is able to balance the vacuum pressure, these $\chi$ remnants form a bound state which is known as Fermi-ball Lee:1986tr ; Hong:2020est ; Kawana:2021tde ; DelGrosso:2023trq . These Fermi-balls formed at $T_{*}$ defined by $p(T_{*})$=0.29. The energy of a Fermi-ball can be written as
@@
$E_{Y}\simeq-\frac{3g_{\chi}^{2}}{8\pi}\frac{Q_{\rm FB}^{2}}{R_{\rm FB}}\left(\frac{L_{\phi}}{R_{\rm FB}}\right)^{2}.$ (13)
@@
$\beta_{\rm PBH}\gtrsim\beta_{\text{crit}}\simeq\frac{T_{\rm evap}}{T_{n}}\,,$ (16)
@@
$T_{n}=1.61\times 10^{18}~{}\beta_{\rm PBH}\left(\frac{100}{g_{*}}\right)^{1/4}\left(\frac{M_{\rm in}}{1\hskip 2.84544pt\rm g}\right)^{-1/2}\left(\frac{\eta_{\chi}}{10^{-3}}\right)^{-1}\frac{1}{\alpha_{*}^{1/4}}\hskip 2.84544pt\text{GeV}\,,$ (17)
$\beta/\textbf{H}_{*}=1.75\times 10^{-2}\alpha^{1/4}_{*}\left(\frac{\eta_{\chi}}{10^{-3}}\right)v_{w}\left(\frac{g_{*}}{100}\right)^{1/12}\beta_{\text{PBH}}^{-2/3}\,.$ (18)
@@ V GW from PBH density fluctuations
While there are several ways in which PBH can give rise to GW Anantua:2008am ; Zagorac:2019ekv ; Saito:2008jc ; Hooper:2020evu , for our purpose of ultralight PBH, the GW spectrum generated by most of these mechanisms lies at very high frequencies. Hence, we rather focus on GW generated because of the inhomogeneity in the PBH distribution after they are formed Papanikolaou:2020qtd ; Domenech:2020ssp ; Domenech:2021wkk ; Domenech:2021ztg ; Papanikolaou:2022chm . Such a spectrum is independent of the formation mechanism of PBH, and is within reach of near-future GW detectors.
@@
where $k_{\rm UV}$ corresponds to the comoving wavenumber corresponding to the mean separation between PBH, whereas $k_{\rm eq}$, $k_{\rm eva}$ indicates the comoving wavenumber of the modes entering the horizon when PBH begins to dominate and when they evaporate respectively. The ratio of the comoving wavenumbers are found using Domenech:2020ssp $k_{\rm UV}/k_{*}=(\frac{3M_{\rm in}}{4\pi\rho_{\rm PBH}^{\rm in}})^{-1/3}\bf{H_{*}}^{-1}$, $k_{\rm eq}/k_{*}=\sqrt{2}\beta_{\rm PBH}^{2/3}$ and $k_{\rm eva}/k_{*}=(\beta_{\rm PBH}H_{\rm eva}/\bf{H_{*}})^{1/3}$, where $\rho_{\rm PBH}^{\rm in}$, $k_{*}$ and $H_{\rm eva}$ ($\bf{H_{*}}$) denotes the initial energy density of PBH, wavenumber at the time of PBH formation, Hubble parameter at the time of PBH evaporation (formation) respectively. Expressing explicitly in terms of the PBH formation temperature $T_{*}\simeq T_{n}$, we get
$\frac{k_{\rm UV}}{k_{\rm eva}}\simeq 10^{6}\left(\frac{M_{\rm in}}{10^{4}\rm g}\right)^{2/3},$ (32)
$\frac{k_{\rm eq}}{k_{\rm UV}}\simeq 3.18\times 10^{-11}(\beta_{\rm PBH})^{2/3}\left(\frac{M_{\rm in}}{1~{}\rm g}\right)^{1/3}\left(\frac{T_{n}}{1~{}\rm GeV}\right)^{2/3}.$ (33)
Using the above ratio of wavenumbers in Eq. (31), we can write the GW amplitude as
@@
Note that for PBH formed through Fermi-ball collapse, $T_{n}$ in turn depends on the PBH parameters and the asymmetry $\eta_{\chi}$ through Eq. (17).
@@
$\beta_{\rm PBH}\lesssim\beta_{\rm max}\simeq 1.8\times 10^{-4}\left(\frac{M_{\rm in}}{10^{4}g}\right)^{-17/32}\left(\frac{\alpha}{0.1}\right)^{3/16}\left(\frac{\beta/{\bf H_{*}}}{10}\right)^{3/4}\,.$ (35)